HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

THE 2,4-CHALLENGE IN THE MICROWAVE SPECTRA OF 2,4-DIMETHYLTHIAZOLE AND 2,4-DIMETHYLFLUOROBENZENE: STRATEGY TO ASSIGN VERY LOW BARRIER METHYL INTERNAL

ROTATION

\underline{S. KHEMISSI} ,^{1} M. SCHWELL,^{1} I. KLEINER,^{2} H. V. L. NGUYEN^{1,3},^{1}Univ Paris Est Creteil and Université Paris Cité, CNRS, LISA, F-94010 Créteil, France, 2Université Paris Cité and Univ Paris Est Creteil, CNRS, LISA, F-75013 Paris,

France,^{3}Institut Universitaire de France (IUF), F-75231 Paris, France

The analysis and modeling of internal rotations with extremely low barriers is a challenging task in microwave spectroscopy. An almost free internal rotation hindered by a barrier less than 20 cm^{−1} causes large torsional splittings in the spectrum. For molecules with two inequivalent methyl groups, the methyl rotations couple with each other, especially in aromatic molecules through π conjugation, and all rotational lines split into quintets, making spectral assignment and modeling very difficult. The spectral analysis becomes even more complicated if additional effects are present such as the ^{14}N nuclear quadrupole coupling, which causes quadrupole hyperfine splittings in addition to the torsional splittings.^{1} Probably for these reasons, a very limited number of studies on aromatic systems undergoing nearly barrier-free internal rotations have been reported.^{2} We will present here the challenges for two such molecules, 2,4-dimethylthiazole (24DMTA) and 2,4-dimethylfluorobenzene (24DMFB). The microwave spectra were recorded using a pulsed molecular jet Fourier transform microwave spectrometer operating in the frequency range from 2.0 to 26.5 GHz.^{3} The experimental work was supported by quantum chemical calculations. For both molecules, the XIAM_{m}_{o}_{d} program^{4}^{,}^{5} was used at the beginning to fit the first assigned lines, but it fails to predict new lines at higher J and K values, as well as to achieve the measurement accuracy of 2 kHz for the final data set.^{6} To assign more lines, fitting each torsional species separately using odd power order parameters to get reliable predictions using the WS18 code^{7} and the SFLAMS code,^{4} as well as checking the assigned frequencies by combination difference loops (Ritz cycles)^{8} is the key to success. Finally, the BELGI-C_{s}-2Tops-hyperfine^{9} and BELGI-C_{s}-2Tops^{10} programs were used for 24DMTA and 24DMFB, respectively, to decrease the standard deviation to measurement accuracy in a global fit. For 24DMTA, the barriers to internal rotation of the 2- and 4-methyl groups were determined to be 19.070(58) cm^{−1} and 396.707(25) cm^{−1}, respectively. They are 3.23(40) cm^{−1} and 227.039(51) cm^{−1} for 24DMFB.

1doi.org/10.1515/psr-2020-0037 , H.V.L. Nguyen, I. Kleiner, Phys. Sci. Rev. 7, 679−726 (2020).

2doi.org/10.3390/molecules27123948 , H.V.L. Nguyen, W. Caminati, J.-U. Grabow, Molecules. 27, 3948 (2022).

3doi.org/10.1063/1.1147553 , J.-U. Grabow, W. Stahl, H. Dreizler, Rev. Sci. Instrum. 67, 4072−4084(1996).

4doi.org/10.1063/1.5142401 , S. Herbers, S.M. Fritz, P. Mishra, H.V.L. Nguyen, T.S. Zwier, J. Chem.

Phys. 152, 074301 (2020).

5doi.org/10.1016/j.jms.2020.111289 , S. Herbers, H.V.L. Nguyen, J. Mol. Spectrosc. 370, 111289 (2020).

6doi.org/10.1515/zna-1990-0817 , J.-U. Grabow, W. Stahl, Z. Naturforsch. 45a, 1043−1044 (1990).

7doi.org/10.1063/5.0049418 , T. Nguyen, W. Stahl, H.V.L. Nguyen, I. Kleiner, J. Chem. Phys. 154, 204304 (2021).

p-number: p160  Submitted on Wed, 14 Jun 2023 23:17:28 +0200

8doi.org/10.1016/j.jms.2015.10.005 , Y. Zhao, H.V.L. Nguyen, W. Stahl, J.T. Hougen, J. Mol. Spectrosc. 318, 91−100 (2015).

9doi.org/10.1063/5.0049418 , V. Van, T. Nguyen, W. Stahl, H.V.L. Nguyen, I. Kleiner, J. Mol. Struct. 1207, 127787 (2020).

10doi.org/10.1016/j.jms.2011.07.005 , M. Tudorie, I. Kleiner, J.T. Hougen, S. Melandri, L.W. Sutikdja, W. Stahl, J. Mol. Spectrosc. 269, 211−225 (2011).

p-number: p160  Submitted on Wed, 14 Jun 2023 23:17:28 +0200